OpenStudy (anonymous):

Better way to express the domain of sec, csc, and cot functions? Lol, my textbook writes it in an ugly and confusing way. As an example, they write domain of sec functions like x does not equal pi/2, +/- 3pi/2.... Is there an easier way to express this? Maybe like x does not equal pi/4+(k*pi/2)? Is this correct?

OpenStudy (anonymous):

Can someone clarify this for me? I have a test tomorrow lol...

myininaya (myininaya):

\[secx=\frac{1}{cosx}\] where is cosx,0? cosx=0 when x=pi/2, 3pi/2, and so on whenever we go around a circle that is 2pi so we can keep going around a circle forever and forever so cosx=0 when x=pi/2+2npi and x=3pi/2+2npi for n=...,-3,-2,-1,0,1,2,3,....

OpenStudy (anonymous):

Domain of Cot and Cosec - all R except integer multiples of pi Sec-all R except odd integer multiples of pi/2

OpenStudy (anonymous):

Sec is same as Tan

myininaya (myininaya):

because tan=sin/cos

myininaya (myininaya):

they have the same denominator

OpenStudy (anonymous):

Sorry, I am still confused ><

myininaya (myininaya):

about?

myininaya (myininaya):

can you narrow down for me what you are confused about?

OpenStudy (anonymous):

The domain of sec x and csc x.. about +2npi

myininaya (myininaya):

n is the number of times we go around a circle

OpenStudy (radha):

domain of y=cot x.......x belongs to R-(n pi) domain of y=cosec x............x belongs to (-infinity,-1]union[1, infinity) domain of y=sec x.............x belongs to R-(2n+1)pi/2 (here n is a natural number)

myininaya (myininaya):

we can go counterclockwise or clockwise that is why we have the negative and the positive n

OpenStudy (anonymous):

Oh no, not about the n, about why you would want to add 2npi in general. Like why can't you add it by the period ?

myininaya (myininaya):

?

OpenStudy (anonymous):

Uhm, I mean like so the sec function has asymptotes right? There is an asymptote on -3pi/2, -pi/2, pi/2, 3pi/2. So I understand how my book lists it like that, but I think it is ugly.

myininaya (myininaya):

\[\cos(\frac{\pi}{2}+2npi)=\cos(\frac{\pi}{2})\cos(2npi)-\sin(\frac{\pi}{2})\sin(2npi)=0(1 or -1)-1(0)=0-0=0\]

myininaya (myininaya):

\[\cos(\frac{3\pi}{2}+2npi)=\cos(\frac{3\pi}{2})\cos(2npi)-\sin(\frac{3\pi}{2})\sin(2npi)=0(-1 or 1)-1(0)=0-0=0\]

myininaya (myininaya):

yes we have vertical asymptotes at \[x=\frac{\pi}{2}+2npi, x=\frac{3\pi}{2}+2npi, n \in \mathbb{Z}\]

OpenStudy (anonymous):

What about in 3csc(3x+pi)-2? So I first factor out the 3 = y = 3csc(3(x+pi/3))-2 I know my period is 2pi/3 So what would be my domain?

myininaya (myininaya):

\[\csc(3x+\pi)=\frac{1}{\sin(3x+\pi)}\] we need to figure out when sin(3x+pi) is 0 so remember sin(k) is zero when \[k=\pi+2npi, k=2\pi+2npi, k \in \mathbb{Z}=> 3x+\pi=2\pi+2npi, 3x+\pi=\pi+2npi\] now we need to solve for x to find what the domain should exclude

myininaya (myininaya):

\[x=\frac{\pi+2npi}{3}, x=\frac{2npi}{3}\]

myininaya (myininaya):

the domain is all real numbers except \[x=\frac{\pi+2npi}{3}, x=\frac{2npi}{3} \]

OpenStudy (anonymous):

How did you get this? "x=2npi/3" ? I understand how you got x=pi + 2npi/3..

myininaya (myininaya):

we have 3x+pi=pi+2npi subtract pi on both sides 3x=2npi divide both sides by 3 x=2npi/3